OpenStudy (anonymous):

Chuck's CheeseHouse makes two types of cheese; mozzarella and parmesan. Mozzarella costs $2.60 per pound and parmesan costs $3.30 per pound. How much parmesan cheese went into making a 60 pound mixture worth $3.10 per pound? Round to the nearest pound.(Points : 1) 17 lbs 23 lbs 43 lbs 37 lbs

OpenStudy (anonymous):

43 lbs Let m and p be the pound weight of mozzarella and parmesan cheese respectively. Solve the following equations for m and p:\[\left\{m+p=60,\frac{26}{10}m+\frac{33 }{10}p=\frac{31}{10}*60\right\}\]\[\left\{m=\frac{120}{7},p=\frac{300}{7}\right\} \]\[\text{Round}\left[\frac{300}{7},1\right]\to 43 \]